3 9 16 48 55 165 172 --- what are the next 3 numbers in this sequence and why
Answer by JulietG(1812)

Here's one solution:
3 *3 = 9 +7 = 16
16 *3 = 48 +7 = 55
55 *3 = 165 +7 = 172
172 *3 = 516 +7 = 523
523 *3 = 1569 +7 = 1576